OpenStudy (anonymous):

1. Mr. Martinez is putting together supply boxes for his students. He has 72 pens, 36 pencils, and 54 tablets to divide among the boxes. He wants to divide the items evenly and not have any items left over. a.) List the common factors of 72, 36, and 54. Which common factor represents the greatest number of supply boxes he can make? Explain. b.) If he makes the greatest number of supply boxes, how many of each item would go in each box? Show how you know.

Directrix (directrix):

Factorization of 72, 36, and 54. 72 = 2^3 * 3^2 36 = 2^2 * 3^2 54 = 3^3 * 2

OpenStudy (anonymous):

Thank you @Directrix